Jeff's bank offers a savings account, compounded monthly, with an APR of
2.25%.
What principal P should be invested so that the account balance is $2500 in three years? (Assume the interest is compounded monthly. Round your answer to two decimal places.)

Answer by stanbon(75887)

Jeff's bank offers a savings account, compounded monthly, with an APR of
2.25%.
What principal P should be invested so that the account balance is $2500 in three years?
(Assume the interest is compounded monthly. Round your answer to two decimal places.)
--------------------
A(t) = P(1+(r/12))^(12t)
----------------------
2500 = P(1+0.0225/12)^(36)
-------
2500 = P(1.0698)
Ans:: P = $2336.97
----------------
Cheers,
Stan H.
----------------------------